# HARec: Hyperbolic Graph-LLM Alignment for Exploration and Exploitation in Recommender Systems

### v1

@@ 5.3. Ablation Study
Figure 4 shows that our full model outperforms all ablation variants, highlighting its overall effectiveness. Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the importance of each component and their effective integration.
5.4. Hierarchical Structure Analysis
To validate the effectiveness of our hierarchical structure, which allows users to control their level of exploration within recommendations, we conducted a series of analyses. Our goal is to demonstrate how the hierarchical model provides a structured balance between exploration and exploitation, allowing users to adjust the generality of recommendations by themselves.
First, we analyze the norms of node embeddings across hierarchical layers. As shown in Table 4, nodes in higher layers have consistently smaller norms. This observation supports our design intention: higher-layer nodes represent broader, generalized group preferences, aligning with findings in Section 4.5. This structure enables more abstract grouping as the layers ascend, thereby capturing progressively broader user interests.
To further examine the hierarchy’s practical utility, we designed experiments that simulate varying user preferences in daily interactions. Using a temperature parameter $\tau=0.2$ to control recommendation randomness, we tested recommendation performance across different hierarchical levels, from layer 00 (original recommendation level) up to layer $3$. The results, visualized in Figure 5, reveal a clear pattern: as the level $l$ increases, utility metrics tend to decrease, while diversity metrics increase.
This progression highlights the intended trade-off our model offers: recommendations from lower layers (e.g., layer 00) focus on utility and specific user interests, ideal for exploitation. In contrast, higher layers (e.g., layer $3$) prioritize diversity by presenting broader, less tailored options, thus encouraging exploration. This systematic change across levels provides users with a flexible mechanism to adjust recommendation focus: from niche, highly relevant items to a more varied, exploratory set—based on their current preference for exploration or exploitation.
To underscore the effectiveness of semantic alignment, particularly for cold-start (tail) items, we conducted experiments measuring utility performance on both head and tail items. In our context, head items (H20) refer to the top $20\%$ of items with the highest number of interactions in the training dataset, representing the most popular items. The remaining items are categorized as tail items (T80), which have limited interaction history, representing less frequently engaged (cold-start) items.
6. CONCLUSION
@@
In this work, we introduce HARec, a novel hyperbolic graph-text alignment framework that effectively balances exploration and exploitation by integrating semantic and collaborative information, with theoretical analysis supporting the advantages of this approach within hyperbolic space. Additionally, we propose a unique mechanism that enables personalized control over the exploration-exploitation trade-off through a hierarchical tree structure, allowing users to adjust recommendation preferences dynamically. Extensive experiments demonstrate that HARec achieves state-of-the-art performance, with up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics, making it the first model to excel in both criteria simultaneously, a significant achievement that has yet to be attained by any previous models.
Appendix A Appendix
In the supplementary materials, we provide additional details on our datasets and present several case studies illustrating the performance of our decoder module.
A.1. Profile Decoder
@@
To further evaluate the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks. Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers. These Euclidean-mapped embeddings are then processed through a mixture of experts (MoE) module. The MoE serves to refine the embeddings, ensuring consistency in both dimensionality and semantic representation (Ma et al., 2024). Once processed, these embeddings are fed into the language model, specifically after the positional layer, to incorporate them within a sequential architecture. To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
Details of the input prompt structure and processing flow are depicted in Figure 6. To facilitate seamless embedding integration, we introduce two additional tokens, ¡USER_EMBED¿ and ¡ITEM_EMBED¿, to the language model’s tokenizer. This addition enables these embeddings to be treated as single, coherent tokens, ensuring accurate tokenization and embedding interpretation. After passing through the positional layer, the placeholders are dynamically replaced by the adapted embeddings generated by the MoE, allowing the embeddings to retain context-specific information.
A.2. Decoder Quality
@@
To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder. The results are shown in Table 6. We use BERTScore (Zhang et al., [n. d.]) to evaluate the semantic alignment of generated sentences with the original profile. Results indicate that our HARec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information. This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
A.2.1. Decoder Case Study
@@
————————————————————————————————————— Original user profile:
@@
————————————————————————————————————— In this example, the profile recovered by our decoder accurately captures the user’s preferences for Thai and Chinese cuisines, broadly summarized as an affinity for Asian fusion flavours. Additionally, our decoder identifies the user’s interest in diverse culinary experiences, indicating a propensity toward exploration.
————————————————————————————————————— Original item profile:
@@
————————————————————————————————————— The alignment between the original and recovered item profiles shows that our decoder effectively captures key thematic elements of Ceol Irish Pub. The original profile emphasizes aspects such as happy hour, a wide selection of drinks, live music, and a welcoming atmosphere with attentive service. The recovered profile distills these characteristics, focusing on the pub’s appeal for users who enjoy diverse beer options, live music, and a relaxed setting. This alignment suggests that our decoder reliably summarizes key attributes, albeit with a degree of abstraction.
Figure 6. A depiction of model prompt instruction.
Table 6. Experiments assessing decoder quality.

### v2

@@ 5.3 Ablation Study
Figure 3 shows that our full model outperforms all ablation variants, highlighting its overall effectiveness. Models lacking both margin ranking and semantic alignment perform the worst, while removing either module results in decreased utility, emphasizing the effectiveness of integration.
5.4 Hierarchical Representation Structure Analysis
We evaluated the effectiveness of our hierarchical representation structure through experiments simulating varying user preferences. Using a temperature parameter $\tau=0.5$ to control recommendation randomness, we analyzed performance across various hierarchical layers. As shown in Figure 4, utility metrics increase while diversity metrics decrease with larger layers. This trade-off aligns with the model’s design: larger layers focus on utility and specific interests for exploitation, while smaller layers prioritize diversity for exploration. This structure provides users with a flexible mechanism to adjust recommendations from niche, relevant items to broader, exploratory sets based on their preferences. A case study is provided in Appendix C.
Additionally, to validate the hierarchical structure’s capability to capture varying degrees of group preferences, we analyzed the average norm of node embeddings across layers (Figure 5). The results reveal a consistent increase at larger layers (farther from the center in hyperbolic space). As previously mentioned, hyperbolic geometry ensures that popular items or exploratory users are positioned closer to the origin, while niche items or users are positioned toward the boundaries. This finding aligns seamlessly with the properties of hyperbolic space, validating the intended design of the hierarchical structure: higher-level (smaller-layer) nodes effectively encapsulate abstract and aggregated preferences, whereas lower-level (larger-layer) nodes focus on fine-grained, personalized preferences.
6 CONCLUSION
@@
In this work, we present HERec, a hyperbolic graph-LLM framework that effectively balances exploration and exploitation in recommender systems. Our approach integrates semantic and collaborative information within hyperbolic space, leveraging the unique properties of hyperbolic geometry. Furthermore, we introduce a personalized mechanism for balancing the exploration-exploitation trade-off using a hierarchical representation structure, empowering users to adjust recommendation preferences flexibly. Extensive experiments demonstrate that our model achieves state-of-the-art performance, achieving up to a $5.49\%$ improvement in utility metrics and an $11.39\%$ increase in diversity metrics. Remarkably, HERec is the first framework to excel in both utility and diversity simultaneously, marking a significant advancement in the field.
ACKNOWLEGDEMENTS
This work was partially supported by the National Science Foundation (NSF) IIS Div Of Information & Intelligent Systems 2403317. We also gratefully acknowledge partial support from an Amazon Research Award, funding generously provided by Snap Research, and resources made available by the Yale Office of the Provost.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proof
In this section, we provide a rigorous mathematical proof of Proposition 4.1, demonstrating the unique advantage of the aligning mechanism in the hyperbolic space.
A.1 Gradient Analysis in Hyperbolic Space.
Suppose $\mathbf{x}$ is a hyperbolic embedding derived from  4.1, and $\mathbf{y}$ is the corresponding semantic embedding from  4.2, the hyperbolic distance is given by:
$d_{\mathcal{H}}(\mathbf{x},\mathbf{y})=\operatorname{arccosh}(z)$ (19)
where $z=-\langle\mathbf{x},\mathbf{y}\rangle_{\mathcal{L}}=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}$.
For large $\|\mathbf{x}\|$ and $\|\mathbf{y}\|$, given $x_{0}=\sqrt{1+\|\mathbf{x}\|^{2}}$ and $y_{0}=\sqrt{1+\|\mathbf{y}\|^{2}}$, we adopt the following approximations for further calculation:
$x_{0}\approx\|\mathbf{x}\|,\quad y_{0}\approx\|\mathbf{y}\|,\quad\sqrt{z^{2}-1}\approx z$ (20)
Thus, the Lorentz product can be approximated as:
$z=x_{0}y_{0}-\sum_{i=1}^{n}x_{i}y_{i}\approx\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)$ (21)
Let $x_{j}=\|\mathbf{x}\|\hat{x}_{j}$ and $y_{j}=\|\mathbf{y}\|\hat{y}_{j}$ (where $\hat{x}_{j}$ represents the direction of $x_{j}$). The partial derivative of $z$ with respect to $x_{j}$ is calculated as follows:
$\frac{\partial z}{\partial x_{j}}$ $=\frac{x_{j}}{x_{0}}y_{0}-y_{j}$ (22)
$\approx\frac{x_{j}}{\|\mathbf{x}\|}\|\mathbf{y}\|-y_{j}$
$=\|\mathbf{y}\|(\hat{x}_{j}-\hat{y}_{j})$
Therefore, the gradient norm of $d_{\mathcal{H}}$ for $\mathbf{x}$ can be approximated by:
$\|\nabla_{\mathbf{x}}d_{\mathcal{H}}\|$ $=\frac{\|\nabla_{\mathbf{x}}z\|}{\sqrt{z^{2}-1}}$ (23)
$\approx\frac{\|\nabla_{\mathbf{x}}z\|}{z}$
$=\frac{\|\mathbf{y}\|\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|\|\mathbf{y}\|(1-\cos\theta)}$
$=\frac{\|\hat{\mathbf{x}}-\hat{\mathbf{y}}\|}{\|\mathbf{x}\|(1-\cos\theta)}$
This result suggests that nodes with large norms experience smaller gradient updates, preserving local structures, while nodes with smaller norms undergo larger updates, refining global hierarchical relationships and effectively capturing hierarchical structures. This aligns with the properties of hyperbolic geometry, where small positional changes near the origin result in significant distance changes, and vice versa. Our dynamic gradient update mechanism is well-suited to this characteristic, enabling precise adjustments and enhancing the accuracy of the updates.
A.2 Gradient Analysis Euclidean Space.
For comparison, we examine the gradient behavior in Euclidean space. The Euclidean distance is defined as:
$f(\mathbf{x})=d_{E}(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|$ (24)
The gradient of $f(\mathbf{x})$ with respect to $\mathbf{x}$ is:
$\nabla_{\mathbf{x}}f(\mathbf{x})=\frac{\mathbf{x}-\mathbf{y}}{\|\mathbf{x}-\mathbf{y}\|}=\frac{\mathbf{x}-\mathbf{y}}{d_{E}(\mathbf{x},\mathbf{y})}$ (25)
which has a magnitude of:
$\|\nabla_{\mathbf{x}}f(\mathbf{x})\|=1$ (26)
The constant gradient magnitude demonstrates Euclidean space’s uniform update behavior, lacking intrinsic adaptation to hierarchical structures. This fundamental difference in gradient dynamics explains hyperbolic space’s superiority for hierarchical representation learning.
Appendix B Profile Decoder
Figure 6: A depiction of model prompt instruction.
@@ Appendix B Profile Decoder
To further evaluate the effectiveness of hyperbolic alignment and the expressiveness of our learned hyperbolic embeddings, we introduce a profile decoder designed to reconstruct the original profile from these embeddings, thereby broadening their application beyond recommendation tasks. Our method begins by mapping the hyperbolic embeddings back into Euclidean space to facilitate alignment with traditional network layers. These Euclidean-mapped embeddings are then processed through a mixture of experts (MoE) module. The MoE serves to refine the embeddings, ensuring consistency in both dimensionality and semantic representation (Ma et al., 2024). Once processed, these embeddings are fed into the language model, specifically after the positional layer, to incorporate them within a sequential architecture. To enhance training efficiency and minimize computational cost, we freeze the language model’s parameters, making the MoE module the only trainable component in the architecture.
[…]
B.1 Decoder Quality
@@
To assess the effectiveness of our decoder mechanism, we compare the embeddings from our model against those from Euclidean embeddings (LightGCN) and Hyperbolic embeddings (HICF) as inputs to the decoder. The results are shown in Table 5. We use BERTScore (Zhang et al., 2019) to evaluate the semantic alignment of generated sentences with the original profile. Results indicate that our HERec outperforms both Euclidean and hyperbolic baselines, demonstrating that our model’s embeddings support recommendation and retain user/item profile information. This finding underscores that our model captures collaborative information alongside recoverable language semantics: a capability not achieved by previous models.
B.2 Decoder Case Study
@@
 Original user profile:
@@
In this example, the profile recovered by our decoder accurately captures the user’s preferences for Thai and Chinese cuisines, broadly summarized as an affinity for Asian fusion flavours. Additionally, our decoder identifies the user’s interest in diverse culinary experiences, indicating a propensity toward exploration.
 Original item profile:
@@
The alignment between the original and recovered item profiles shows that our decoder effectively captures key thematic elements of Ceol Irish Pub. The original profile emphasizes aspects such as happy hour, a wide selection of drinks, live music, and a welcoming atmosphere with attentive service. The recovered profile distills these characteristics, focusing on the pub’s appeal for users who enjoy diverse beer options, live music, and a relaxed setting. This alignment suggests that our decoder reliably summarizes key attributes, albeit with a degree of abstraction.
Table 5: Experiments assessing decoder quality.
@@
Appendix C Hierarchical Structure Case Study
In this section, we present a case study examining the exploration-exploitation trade-off in the hyperbolic hierarchical structure. We randomly select a user and identify items that share different layers of the lowest common ancestor (LCA) with the user in the hierarchy tree. Ideally, a larger LCA layer indicates that the user and item were grouped earlier in the hierarchy, suggesting a stronger likelihood of interaction. Conversely, a smaller LCA layer implies that the user and item were grouped only after multiple iterations, indicating a weaker likelihood of interaction.
 User Profile:
The user enjoys books with supernatural elements, horror and paranormal romance. They also appreciate fast-paced action and suspenseful thrillers with political intrigue and espionage. The user likes complex characters dealing with dark themes.
Item 1 Profile (LCA=10):
The book ’Spirit Fighter (Son of Angels, Jonah Stone)’ would likely appeal to young readers who enjoy stories about supernatural beings and the fight between good and evil. It is a book that offers entertainment and inspiration, as well as a glimpse into the supernatural world.
Item 2 Profile (LCA=5):
Fans of contemporary romance novels will enjoy Barefoot Summer (A Chapel Springs Romance). This book combines heartwarming romance, relatable heartaches, and lighthearted humor to create a delightful and engaging story. It is ideal for those looking for an easy-to-read and entertaining summer novel.
As demonstrated in the case study, the user shows a preference for supernatural elements, which aligns closely with the description of item 1, indicating a high likelihood of interaction. In contrast, while the user profile mentions a preference for dark themes, item 2 focuses on romance and heartwarming narratives, suggesting a low probability of user interest in item 2.
This analysis aligns seamlessly with the intended design of our hierarchical tree, further validating the effectiveness of our model in capturing and leveraging nuanced user-item relationships.